Project

General

Profile

Actions

MOM update 20170926

  • Absorber full last night
  • Tracker Cryo 2 PSU instability: dodgy fuse connection
  • DS IOC issue: last problem on Saturday
  • Status of alarms: new alarms, since the severity level was not implemented by default for a lot of alarms. First iteration with experts has to be implemented.
  • Open shifts
  • Summary of data taken
    Tag Time(H:M) Dips P.Trig(k) @ TOF1 TOF2 Runs
    ToF-Calib+150 4:16 12039 528 101 9628,9629,9630
    PionReference-NoDS 0:55 2598 132 42 9610,9612
    10-140+M3-Test3 1:32 4345 459 108 9664
    ToF-Calib-NoDS 2:22 6667 633 130 9618
    Individual Settings 3:41 4256 265 146 9453,9454,9455,9609,9641,9645,9653,9657
    ToF-Calib+150-NoDS 7:32 21114 736 151 9621,9622,9623,9625,9626
    400MeV+DS200 2:50 7932 933 360 9619,9620
    3-140+M3-Test3 16:34 45242 3111 2556 9654,9659,9660,9661,9677,9678,9679,9683,9687,9690,9691
    6-140+M3-Test2 13:41 38505 3198 1998 9663,9665,9666,9671,9676,9680,9681,9684,9688
    10-140+M3-Test4 10:29 29503 3639 1006 9667,9668,9670,9672,9673,9674,9685,9686,9689
  • 64 hours (out of 164 hours since 8am 19/9)
  • TOF calibration
  • Detectors alignment
  • Beam line tuning
  • 2017-02-1 with "empty" LH2 absorber
  • 2017-02-1 with full LH2 absorber (yesterday 10pm)
    Tag P.Trig(k) @ TOF1
    3-140+M3-Test3 1170
    6-140+M3-Test2 1134
    10-140+M3-Test4 1292
  • Run plan:
    • Tomorrow :change currents to 2017-02-2 for the 200 MeV/c beams
    • Friday: move to flip mode in 2017-02-3 setting

Updated by Franchini, Paolo about 6 years ago · 6 revisions